COUPLED-CHANNEL ANALYSIS OF THE D^{1}Π − d^{3}Π COMPLEX IN NaK; POTENTIAL ENERGY CURVES AND SPIN-ORBIT FUNCTIONS

ANASTASIA DROZDOVA and \underline{AMANDA J. ROSS}, LASIM, Université Lyon 1 & CNRS, 43 Bd du 11 novembre 1918, F-69622 Villeurbanne, France ; ANDREY V. STOLYAROV, Department of Chemistry, Moscow State University, 119992 Moscow, Russia ; WLODZIMIERZ JASTRZȨBSKI, Institute of Physics, Polish Academy of Sciences, Al. Lotnikóv 32/46, 02-668 Warsaw, Poland ; and PAWEŁ KOWALCZYK, Institute of Experimental Physics, Warsaw University, ul. Hoża 69, 00-681 Warsaw, Poland.

Two-colour polarization labeling experiments measuring the D ← X system of NaK have characterised more than 99 % of the potential well of the D ^{1}Π state of NaK^{a}, the last observed level being located 7 cm^{−1} below the Na(3p ^{2}P_{3/2}) + K(4s) atomic asymptote. The vibrational progressions all exhibit irregular intervals, because of strong interactions with the nearby d ^{3}Π state. A comprehensive analysis has now been made of all available data concerning the D ^{1}Π and d ^{3}Π^{b} states. The potential curves are represented by Morse/Lennard Jones analytical functions, with ab initio constraints on the long-range part of V(R). Morse functions are also used to represent the R-dependent diagonal and off-diagonal spin-orbit terms. Initial values for the spin-orbit coupling matrix elements were extracted from quasi-relativistic ab initio calculations. In total, 29 parameters were required to recalculate 95 % of the 1400 observed term energies to within experimental uncertainty, giving an unweighted standard deviation 0.03 cm^{−1}.

aA Adohi-Krou et al. J. Mol. Spectrosc. \underline{250} 27 (2008)

bP. Kowalczyk J. Mol. Spectrosc. \underline{136} 1 (1989)